OpenStudy (zaxoanl):

parametric curve x=3t+27/t y=3t-27/t t=-3 find the dy/dx in terms of t and evaluate it at the given value of t

OpenStudy (jiteshmeghwal9):

\[\Large \frac{dy}{dx}={\frac{dy}{dt} \over \frac{dx}{dt}}\]

OpenStudy (jiteshmeghwal9):

\[\Large{\frac{dy}{dx}={\frac{d}{dt}\left( \frac{3t-27}{t} \right) \over \frac{d}{dt}\left( \frac{3t+27}{t} \right)}}\]

OpenStudy (jiteshmeghwal9):

solve it now

OpenStudy (zaxoanl):

i multiply top and bot with 3t-27 and stuck on \[\frac{ t^2+81 }{ t^2-81 }\]

OpenStudy (jiteshmeghwal9):

why did u multiply top & bottom with 3t-27 ?

OpenStudy (zaxoanl):

\[\frac{ \frac{ 3t-27 }{t } }{ \frac{ 3t+27 }{ t } }\] doesnt the t cancel out

OpenStudy (jiteshmeghwal9):

hey it is differentiation not division

OpenStudy (jiteshmeghwal9):

first find dy/dt then find dx/dt finally answer will be\[\Large {\frac{dy}{dt}\over \frac{dx}{dt}}\]